OpenStudy (anonymous):

estimate the answer of 30% of 40 and explain how you got your answer.

OpenStudy (anonymous):

.3*40=12 or 30/100*(40)=1200/100=12 ok?

OpenStudy (anonymous):

Or, once you get the hang of it, just multiply 4 by 3 because you can see that "missing out" the two zeros is the same thing as dividing by 100.